# A Review of Large Language Models and Autonomous Agents in Chemistry

### v1

@@ Abstract
Large language models (LLMs) are emerging as a powerful tool in chemistry across multiple domains. In chemistry, LLMs are able to accurately predict properties, design new molecules, optimize synthesis pathways, and accelerate drug and material discovery. A core emerging idea is combining LLMs with chemistry-specific tools like synthesis planners and databases, leading to so-called “agents.” This review covers LLMs’ recent history, current capabilities, design, challenges specific to chemistry, and future directions. Particular attention is given to agents and their emergence as a cross-chemistry paradigm. Agents have proven effective in diverse domains of chemistry, but challenges remain. It is unclear if creating domain-specific versus generalist agents and developing autonomous pipelines versus "co-pilot" systems will accelerate chemistry. An emerging direction is the development of multi-agent systems using a human-in-the-loop approach. Due to the incredibly fast development of this field, a repository has been built to keep track of the latest studies: https://github.com/ur-whitelab/LLMs-in-science.
Keywords
@@ 1 Introduction
The integration of Machine Learning (ML) and Artificial Intelligence (AI) into chemistry has spanned several decades1, 2, 3, 4, 5, 6, 7, 8, 9, 10. Although applications of computational methods in quantum chemistry and molecular modeling from the 1950s-1970s were not considered AI, they laid the groundwork. Subsequently in the 1980s expert systems like DENDRAL11, 12 were expanded to infer molecular structures from mass spectrometry data.13 At the same time, Quantitative Structure-Activity Relationship (QSAR) Models were developed5 that would use statistical methods to predict the effects of chemical structure on activity.14, 15, 16, 17 In the 1990s, neural networks were introduced to domains such as drug design, as summarized well by Yang et al. 5 and Goldman and Walters 18, although they were limited by the computational resources of the time. With an explosion of data from High-Throughput Screening (HTS)19, 20, models then started to benefit from vast datasets of molecular structures and their biological activities. Furthermore, ML algorithms such as Support Vector Machines and Random Forests became popular for classification and regression tasks in cheminformatics,1 offering improved performance over traditional statistical methods.21
@@
Deep learning transformed the landscape of ML in chemistry and materials science in the 2010s.22 Recurrent Neural Networks (RNNs),23, 24, 25, 26, 27 Convolutional Neural Networks (CNNs)28, 29, 30 and later, Graph Neural Networks (GNNs),31, 32, 33, 34, 35, 36 made great gains in their application to molecular property prediction, drug discovery,37 and synthesis prediction.38 Such methods were able to capture complex patterns in data, and therefore enabled the identification of novel materials for high-impact needs such as energy storage and conversion.39, 40
@@
In this review, we explore the next phase of AI in chemistry, namely the use of Large Language Models (LLMs) and autonomous agents. Inspired by successes in natural language processing (NLP), LLMs were adapted for chemical language (e.g., SMILES41) to tackle tasks from synthesis prediction to molecule generation42, 43, 44. We will then explore the integration of LLMs into autonomous agents as illustrated by M. Bran et al. 45 and Boiko et al. 46, which may be used for data interpretation or, for example, to experiment with robotic systems. We are at a crossroads where AI enables chemists to solve major global problems faster and streamline routine lab tasks. This enables, for instance, the development of larger, consistent experimental datasets and shorter lead times for drug and material commercialization. As such, language has been the preferred mechanism for describing and disseminating research results and protocols in chemistry for hundreds of years.47
@@
The review structure is as follows: Section 1.1 discusses current challenges that need to be addressed to increase the impact of AI in chemistry. Section 2 introduces transformers, covering encoder-only, decoder-only, and encoder-decoder architectures. Section 3 surveys work with LLMs, linking each transformer architecture to suitable chemistry areas. Section 4 describes autonomous agents and their applications in chemistry research. Section 5 surveys LLM-based agents in chemistry. Section 6 discusses future challenges and opportunities, followed by the conclusion in Section 7. We distinguish between “text-based” and “mol-based” inputs and outputs, with “text” referring to natural language and “mol” referring to the chemical syntax for material structures, as introduced by Zhang et al. 48.
@@ 1.1 Challenges in chemistry
With the potential to accelerate scientific discovery through the use of AI, it makes sense to connect this potential to some of the more pressing opportunities in the chemical sciences as may be categorized under three broad umbrellas: Property Prediction, Property-Directed Molecule Generation, and Synthesis Prediction. The first challenge is to predict a property for a given compound to decide if it should be synthesized for a specific application, such as an indicator,49 light harvester,50 or catalyst.51 To have better models for property prediction, good quality data is crucial. We discuss the caveats and issues with the current datasets in Section 3.1 and illustrate state-of-the-art findings in Section 3.2.
The second challenge is to generate novel chemical structures that meet desired chemical profiles or specific properties.52 Success would accelerate progress in various chemical applications, but reliable reverse engineering (inverse design)53 is not yet feasible over the vast chemical space.54 For instance, inverse design coupled with automatic selection of novel structures (de novo molecular design) could develop a drug targeting a specific protein while retaining properties like solubility, toxicity, and blood-brain barrier permeability.55 The complexity of connecting de novo design with property prediction is high and raises ethical questions that need to be addressed to guide the development of such models. We show how state-of-the-art models currently perform in Section 3.3.
The third major challenge is predicting its optimal synthesis using inexpensive, readily available, and non-toxic starting materials. There will always likely be a solution within a large chemical space: an alternative molecule with similar properties that is easier to synthesize. Exploring this space to find a new molecule with the right properties and a high-yield synthesis route unites these challenges. The number of possible stable chemicals is estimated to be up to $10^{180}$.56, 57, 58, 59 Exploring this vast space requires significant acceleration beyond current methods.60 Restrepo 56 emphasizes the need to catalog data on failed syntheses to build a comprehensive dataset of chemical features. Autonomous chemical resources can accelerate database growth and tackle this challenge. Thus, automation is considered a fourth major challenge in chemistry.61, 62, 63, 64 The following discussion explores how LLMs and autonomous agents can provide the most value. Relevant papers are discussed in Section 3.4
@@ 2 Large Language Models
The prior state-of-the-art for seq2seq was considered to be the Recurrent Neural Network (RNN),65 typically as implemented in Hochreiter and Schmidhuber 66. The RNN retains “memory” of previous steps in a sequence to predict later parts. However, as sequence length increases, gradients can become vanishingly small or explosively large67, 68, preventing effective use of earlier information in long sequences. RNNs have thus fallen behind Large Language Models (LLMs).
LLMs are deep neural networks (NN) with billions of parameters. Mostly, LLMs implement a transformers architecture, introduced by Vaswani et al. 69, although other architectures for longer input sequences [^1] These longer sequence models are more important in biology, where very long genetic sequences are relevant are being actively explored70, 71, 72. Transformers are well-developed in chemistry and are the dominant paradigm behind nearly all state-of-the-art sequence modeling results and, thus, the focus. More detailed discussion on current LLMs applied to general purposes can be found elsewhere.73
@@ 2.1 The Transformer
The transformer was introduced in, “Attention is all you need” by Vaswani et al. 69 in 2017. A careful line-by-line review of the model can be found in “The Annotated Transformer” 74. The transformer was the first sequence-to-sequence (seq2seq) model based entirely on attention. Here, a seq2seq model processes a text input, such as an English paragraph, and produces a different text sequence, such as a French translation of that paragraph. The concept of “attention” is a focus applied to certain words of the input, which would convey the most importance, or the context of the passage, and thereby would allow for better decision-making and greater accuracy. However, in a practical sense, “attention” is implemented simply as the dot-product between token embeddings with some non-linear learned function, and is described further below.
The transformer architecture implements two main modules: the encoder and the decoder. Figure 1a provides a cartoon diagram of the general encoder-decoder transformer architecture. The input is first tokenized, from the model’s vocabulary,75, 76, 77, 78, 79 into computable integers that are then converted into numerical vectors using embedding layers. A single encoder stack is shown here, where in reality there are many, with six in the original model.69 The later stacks each use outputs from the prior stack, typically as a $\sim$512 per-token vector. Each encoder applies positional encoding typically using sine and cosine functions, with a frequency depending on that word’s position, to handle sequences of any length. Other functions can also be used.80 The encoder stack then comprises a multi-headed self-attention mechanism, which relates each “word” to others in the sequence by computing attention scores based on queries, keys, and values. This is followed by normalization and residual connections. Residual connection is represented in Figure 1a by the “by-passing” arrows, which indicate how each sub-layer output is added to its input, which helps mitigate the vanishing gradient problem.67, 68 Lastly, the output is further refined through a pointwise feed-forward network with an activation function, resulting in a set of vectors representing the input sequence with rich contextual understanding, ready for the decoder.
@@ 2.2 Model training
The common lifetime of an LLM consists of being first pretrained using unsupervised or self-supervised techniques, generating what is called a base model. This base model is then fine-tuned for specific applications using supervised techniques. Finally, the supervised fine-tuned model is further tuned with reward models to improve human preference or some other non-differentiable and sparse desired character81.
@@
A key task after model training is aligning its output with what human users prefer. This can be done for improving LLM style and tone. Or, it can be to reduce harmful outputs because the training objectives used during the pretraining and the fine-tuning do not include human values.82 Reinforcement learning (RL) techniques are among the most common alignment techniques. RL is a type of machine learning where actions are performed and feedback is received through rewards, the goal being to learn a policy that maximizes the cumulative reward. However, human preferences can only be evaluated on complete outputs from models, preventing supervised fine-tuning (which requires token-by-token losses).83, 84, 85 To apply RL, sampling tokens from the model is recast as a Markov Decision Process (MDP) - actions are tokens and the reward is zero until the EOS (end of sequence) token, at which point the human preference reward is applied86 Reinforcement Learning with Human Feedback (RLHF)87 is the most used alignment technique. However, alternative strategies include reinforcement learning with synthetic feedback (RLSF)88, proximal policy optimization (PPO)89, REINFORCE90, and others.83, 86, 91, 92. A general discussion about RL potential for LLM fine-tuning was published by Cao et al. 93 and Shen et al. 82.
There are ways to reformulate the RLHF process into a direct optimization problem with a different loss. This is known as direct preference optimization (DPO)94 and a popular competitor to PPO (the most commonly used approach) due to its simplicity. It overcomes the lack of token-by-token loss signal by comparing two completions at a time. The discussions about which technique is superior remain very active in the literature.95
Finally, the alignment may not be to human preferences but to downstream tasks that do not provide token-by-token rewards. For example, Bou et al. 96 and Hayes et al. 97 both use RL on a language model for improving its outputs on a downstream scientific task.
@@ 2.3 Model types
While the Vaswani Transformer69 employed an encoder-decoder structure for sequence-to-sequence tasks, the encoder and decoder were ultimately seen as independent models, leading to “encoder-only”, and “decoder-only” models described below.
@@ 2.3.1 Encoder-only models
Beyond Vaswani’s transformer,69 used for sequence-to-sequence tasks, another significant evolutionary step forward came in the guise of the Bidirectional Encoder Representations from Transformers, or “BERT”, described in October 2018 by Devlin et al. 98 BERT utilized only the encoder component, achieving state-of-the-art performance on sentence-level and token-level tasks, outperforming prior task-specific architectures.98 The key difference was BERT’s bidirectional transformer pretraining on unlabeled text, meaning the model looks at the context both to the left and right of the word in question, facilitated by a Masked Language Model (MLM). This allows a richer understanding of the sentence. The encoder-only design meant BERT focused on generating a deeper "understanding" of the input sequence, rather than mapping input sequences to output sequences. In pretraining, BERT also uses Next Sentence Prediction (NSP). “Sentence” here means an arbitrary span of contiguous text. The MLM task randomly masks tokens and predicts them by looking at preceding and following contexts simultaneously, inspired by Taylor99. NSP predicts whether one sentence logically follows another, training the model to understand sentence relationships. This bidirectional approach allows BERT to recognize greater nuance and richness in the input data.
Subsequent evolutions of BERT include, for example, RoBERTa, (Robustly optimized BERT approach), described in 2019 by Facebook AI’s Liu et al. 100. RoBERTa was trained on a larger corpus, for more iterations, with larger mini-batches, and longer sequences, improving model understanding and generalization. By removing the NSP task and focusing on the MLM task, performance improved. RoBERTa dynamically changed masked positions during training and used different hyperparameters. Evolutions of BERT also include domain-specific pretraining and creating specialist LLMs for fields like chemistry, as described below (see Section 3).
@@ 2.3.2 Decoder-only models
In June 2018, Radford et al. 101 from OpenAI proposed the Generative Pretrained Transformer (GPT) in their paper, “Improving Language Understanding by Generative Pretraining”. GPT used a decoder-only, left-to-right unidirectional language model to predict the next word in a sequence based on previous words, without an encoder. Unlike the Vaswani Transformer’s decoder, GPT could predict the next sequence, applying general language understanding to specific tasks with smaller annotated datasets.
GPT employed positional encodings to maintain word order in its predictions. Unlike the Vaswani Transformer, GPT’s self-attention mechanism prevented tokens from attending to future tokens, ensuring each word prediction depended only on preceding words. Hence a decoder-only architecture represents a causal language model, one that generates each item in a sequence based on the previous items; the generation of each subsequent output is causally linked to the history of generated outputs and nothing ahead of the current word affects its generation.

### v2

@@ Abstract
Large language models (LLMs) have emerged as powerful tools in chemistry, significantly impacting molecule design, property prediction, and synthesis optimization. This review highlights LLM capabilities in these domains and their potential to accelerate scientific discovery through automation. We also review LLM-based autonomous agents: LLMs with a broader set of tools to interact with their surrounding environment. These agents perform diverse tasks such as paper scraping, interfacing with automated laboratories, and synthesis planning. As agents are an emerging topic, we extend the scope of our review of agents beyond chemistry and discuss across any scientific domains. This review covers the recent history, current capabilities, and design of LLMs and autonomous agents, addressing specific challenges, opportunities, and future directions in chemistry. Key challenges include data quality and integration, model interpretability, and the need for standard benchmarks, while future directions point towards more sophisticated multi-modal agents and enhanced collaboration between agents and experimental methods. Due to the quick pace of this field, a repository has been built to keep track of the latest studies: https://github.com/ur-whitelab/LLMs-in-science.
Keywords Large Language Model, LLM, LLM agent, agent, science, chemistry
@@ 1 Introduction
The integration of Machine Learning (ML) and Artificial Intelligence (AI) into chemistry has spanned several decades1, 2, 3, 4, 5, 6, 7, 8, 9, 10. Although applications of computational methods in quantum chemistry and molecular modeling from the 1950s-1970s were not considered AI, they laid the groundwork. Subsequently in the 1980s expert systems like DENDRAL11, 12 were expanded to infer molecular structures from mass spectrometry data.13 At the same time, Quantitative Structure-Activity Relationship (QSAR) Models were developed5 that would use statistical methods to predict the effects of chemical structure on activity.14, 15, 16, 17 In the 1990s, neural networks, and associated Kohonen Self-Organizing Maps were introduced to domains such as drug design,18, 19 as summarized well by Yang et al. 5 and Goldman and Walters 20, although they were limited by the computational resources of the time. With an explosion of data from High-Throughput Screening (HTS)21, 22, models then started to benefit from vast datasets of molecular structures and their biological activities. Furthermore, ML algorithms such as Support Vector Machines and Random Forests became popular for classification and regression tasks in cheminformatics,1 offering improved performance over traditional statistical methods.23
@@
Deep learning transformed the landscape of ML in chemistry and materials science in the 2010s.24 Recurrent Neural Networks (RNNs),25, 26, 27, 28, 29 Convolutional Neural Networks (CNNs)30, 31, 32 and later, Graph Neural Networks (GNNs),33, 34, 35, 36, 37, 38 made great gains in their application to molecular property prediction, drug discovery,39 and synthesis prediction.40 Such methods were able to capture complex patterns in data, and therefore enabled the identification of novel materials for high-impact needs such as energy storage and conversion.41, 42
@@
In this review, we explore the next phase of AI in chemistry, namely the use of Large Language Models (LLMs) and autonomous agents. Inspired by successes in natural language processing (NLP), LLMs were adapted for chemical language (e.g., SMILES43) to tackle tasks from synthesis prediction to molecule generation44, 45, 46. We will then explore the integration of LLMs into autonomous agents as illustrated by M. Bran et al. 47 and Boiko et al. 48, which may be used for data interpretation or, for example, to experiment with robotic systems. We are at a crossroads where AI enables chemists to solve major global problems faster and streamline routine lab tasks. This enables, for instance, the development of larger, consistent experimental datasets and shorter lead times for drug and material commercialization. As such, language has been the preferred mechanism for describing and disseminating research results and protocols in chemistry for hundreds of years.49
@@
The review structure is as follows: Section 1.1 discusses current challenges that need to be addressed to increase the impact of AI in chemistry. Section 2 introduces transformers, covering encoder-only, decoder-only, and encoder-decoder architectures. Section 3 surveys work with LLMs, linking each transformer architecture to suitable chemistry areas. Section 4 describes autonomous agents and their applications in chemistry research. Section 5 surveys LLM-based agents in chemistry. Section 6 discusses future challenges and opportunities, followed by the conclusion in Section 7. We distinguish between “text-based” and “mol-based” inputs and outputs, with “text” referring to natural language and “mol” referring to the chemical syntax for material structures, as introduced by Zhang et al. 50.
@@ 1.1 Challenges in chemistry
With the potential to accelerate scientific discovery through the use of AI, it makes sense to connect this potential to some of the more pressing opportunities in the chemical sciences as may be categorized under three broad umbrellas: Property Prediction, Property-Directed Molecule Generation, and Synthesis Prediction. The first challenge is to predict a property for a given compound to decide if it should be synthesized for a specific application, such as an indicator,51 light harvester,52 or catalyst.53 To have better models for property prediction, good quality data is crucial. We discuss the caveats and issues with the current datasets in Section 3.1 and illustrate state-of-the-art findings in Section 3.2.
The second challenge is to generate novel chemical structures that meet desired chemical profiles or specific properties.54 Success would accelerate progress in various chemical applications, but reliable reverse engineering (inverse design)55 is not yet feasible over the vast chemical space.56 For instance, inverse design coupled with automatic selection of novel structures (de novo molecular design) could develop a drug targeting a specific protein while retaining properties like solubility, toxicity, and blood-brain barrier permeability.57 The complexity of connecting de novo design with property prediction is high and raises ethical questions that need to be addressed to guide the development of such models. We show how state-of-the-art models currently perform in Section 3.3.
The third major challenge is predicting its optimal synthesis using inexpensive, readily available, and non-toxic starting materials. There will always likely be a solution within a large chemical space: an alternative molecule with similar properties that is easier to synthesize. Exploring this space to find a new molecule with the right properties and a high-yield synthesis route unites these challenges. The number of possible stable chemicals is estimated to be up to $10^{180}$.58, 59, 60, 61 Exploring this vast space requires significant acceleration beyond current methods.62 Restrepo 58 emphasizes the need to catalog data on failed syntheses to build a comprehensive dataset of chemical features. Autonomous chemical resources can accelerate database growth and tackle this challenge. Thus, automation is considered a fourth major challenge in chemistry.63, 64, 65, 66 The following discussion explores how LLMs and autonomous agents can provide the most value. Relevant papers are discussed in Section 3.4
@@ 2 Large Language Models
The prior state-of-the-art for seq2seq had been the Recurrent Neural Network (RNN),67 typically as implemented in Hochreiter and Schmidhuber 68. The RNN retains “memory” of previous steps in a sequence to predict later parts. However, as sequence length increases, gradients can become vanishingly small or explosively large69, 70, preventing effective use of earlier information in long sequences. RNNs have thus fallen behind Large Language Models (LLMs).
LLMs are deep neural networks (NN) with billions of parameters. Mostly, LLMs implement a transformer architecture, introduced by Vaswani et al. 71, although other architectures for longer input sequences [^1] These longer sequence models are more important in biology, where very long genetic sequences are relevant are being actively explored72, 73, 74, 75. Transformers are well-developed in chemistry and are the dominant paradigm behind nearly all state-of-the-art sequence modeling results and, thus, the focus. More detailed discussion on current LLMs applied to general purposes can be found elsewhere.76
@@ 2.1 The Transformer
The transformer was introduced in, “Attention is all you need” by Vaswani et al. 71 in 2017. A careful line-by-line review of the model can be found in “The Annotated Transformer” 77. The transformer was the first sequence-to-sequence (seq2seq) model based entirely on attention, although attention had been a feature for Recurrent Neural Networks (RNNs) some years prior.78 Here, a seq2seq model processes a text input, such as an English paragraph, and produces a different text sequence, such as a French translation of that paragraph. The concept of “attention” is a focus applied to certain words of the input, which would convey the most importance, or the context of the passage, and thereby would allow for better decision-making and greater accuracy. However, in a practical sense, “attention” is implemented simply as the dot-product between token embeddings with some non-linear learned function, and is described further below.
The transformer architecture implements two main modules: the encoder and the decoder. Figure 1a provides a cartoon diagram of the general encoder-decoder transformer architecture. The input is first tokenized, from the model’s vocabulary,79, 80, 81, 82, 83 into computable integers that are then converted into numerical vectors using embedding layers. A single encoder stack is shown here, where in reality there are many, with six in the original model.71 The later stacks each use outputs from the prior stack, typically as a $\sim$512 per-token vector. Each encoder applies positional encoding typically using sine and cosine functions, with a frequency depending on that word’s position, to handle sequences of any length. Other functions can also be used.84 The encoder stack then comprises a multi-headed self-attention mechanism, which relates each “word” to others in the sequence by computing attention scores based on queries, keys, and values. This is followed by normalization and residual connections. Residual connection is represented in Figure 1a by the “by-passing” arrows, which indicate how each sub-layer output is added to its input, which helps mitigate the vanishing gradient problem.69, 70 Lastly, the output is further refined through a pointwise feed-forward network with an activation function (such as ReLU,85 SwiGLU,86 GELU,87 etc) resulting in a set of vectors representing the input sequence with rich contextual understanding, ready for the decoder.
@@ 2.2 Model training
The common lifetime of an LLM consists of being first pretrained using unsupervised or self-supervised techniques, generating what is called a base model. This base model is then fine-tuned for specific applications using supervised techniques. Finally, the supervised fine-tuned model is further tuned with reward models to improve human preference or some other non-differentiable and sparse desired character88.
@@
A key task after model training is aligning the output with what human users prefer. This can be done so as to improve the LLM output’s style and tone, or to reduce harmful outputs because the training objectives used during the pretraining and the fine-tuning may not include human values.89 Reinforcement learning (RL) techniques are among the most common alignment techniques. RL is a type of machine learning where actions are performed and feedback is received through rewards, the goal being to learn a policy that maximizes the cumulative reward. However, human preferences can only be evaluated on complete outputs from models, preventing supervised fine-tuning (which requires token-by-token losses).90, 91, 92 To apply RL, sampling tokens from the model is recast as a Markov Decision Process (MDP) - actions are tokens and the reward is zero until the EOS (end of sequence) token, at which point the human preference reward is applied93 Reinforcement Learning with Human Feedback (RLHF)94 is the most used alignment technique. However, alternative strategies include reinforcement learning with synthetic feedback (RLSF)95, proximal policy optimization (PPO)96, REINFORCE97, and others.90, 93, 98, 99. A general discussion about RL potential for LLM fine-tuning was published by Cao et al. 100 and Shen et al. 89.
There are ways to reformulate the RLHF process into a direct optimization problem with a different loss. This is known as direct preference optimization (DPO)101 and a popular competitor to PPO (the most commonly used approach) due to its simplicity. It overcomes the lack of token-by-token loss signal by comparing two completions at a time. The discussions about which technique is superior remain very active in the literature.102
Finally, the alignment may not be to human preferences but to downstream tasks that do not provide token-by-token rewards. For example, Bou et al. 103 and Hayes et al. 104 both use RL on a language model for improving its outputs on a downstream scientific task.
@@ 2.3 Model types
While the Vaswani Transformer71 employed an encoder-decoder structure for sequence-to-sequence tasks, the encoder and decoder were ultimately seen as independent models, leading to “encoder-only”, and “decoder-only” models described below.
@@ 2.3.1 Encoder-only models
Beyond Vaswani’s transformer,71 used for sequence-to-sequence tasks, another significant evolutionary step forward came in the guise of the Bidirectional Encoder Representations from Transformers, or “BERT”, described in October 2018 by Devlin et al. 105 BERT utilized only the encoder component, achieving state-of-the-art performance on sentence-level and token-level tasks, outperforming prior task-specific architectures.105 The key difference was BERT’s bidirectional transformer pretraining on unlabeled text, meaning the model looks at the context both to the left and right of the word in question, facilitated by a Masked Language Model (MLM). This encoder-only design meant BERT focused on generating a deeper "understanding" of the input sequence, rather than mapping input sequences to output sequences. In pretraining, BERT also uses Next Sentence Prediction (NSP). “Sentence” here means an arbitrary span of contiguous text. The MLM task randomly masks tokens and predicts them by looking at preceding and following contexts simultaneously, inspired by Taylor106. NSP predicts whether one sentence logically follows another, training the model to understand sentence relationships. This bidirectional approach allows BERT to recognize greater nuance and richness in the input data.
Subsequent evolutions of BERT include, for example, RoBERTa, (Robustly optimized BERT approach), described in 2019 by Facebook AI’s Liu et al. 107. RoBERTa was trained on a larger corpus, for more iterations, with larger mini-batches, and longer sequences, improving model understanding and generalization. By removing the NSP task and focusing on the MLM task, performance improved. RoBERTa dynamically changed masked positions during training and used different hyperparameters. Evolutions of BERT also include domain-specific pretraining and creating specialist LLMs for fields like chemistry, as described below (see Section 3).
@@ 2.3.2 Decoder-only models
In June 2018, Radford et al. 108 from OpenAI proposed the Generative Pretrained Transformer (GPT) in their paper, “Improving Language Understanding by Generative Pretraining”. GPT used a decoder-only, left-to-right unidirectional language model to predict the next word in a sequence based on previous words, without an encoder. Unlike the Vaswani Transformer’s decoder, GPT could predict the next sequence, applying general language understanding to specific tasks with smaller annotated datasets.
GPT employed positional encodings to maintain word order in its predictions. Unlike the Vaswani Transformer, GPT’s self-attention mechanism prevented tokens from attending to future tokens, ensuring each word prediction depended only on preceding words. Hence a decoder-only architecture represents a so-called causal language model, one that generates each item in a sequence based on the previous items; the generation of each subsequent output is causally linked to the history of generated outputs and nothing ahead of the current word affects its generation.